OpenStudy (anonymous):

i am doing homework and i am doing measurements in triangles and i ave gotten as far as 2 8/12x = 65 22/12 and i am stuck on how to go further please help

OpenStudy (vijay):

can u write again with brackets ?

OpenStudy (anonymous):

(2 8/12)x = (65 22/12) is this what you mean?

OpenStudy (vijay):

(65 22/12) means ? 65(22/12)

OpenStudy (anonymous):

yes both numbers are fractions and i have to find the value of x. and i am not sure how to simplify them (if they need to be) and i dont know how to divide them i hate fractions!!

OpenStudy (vijay):

still, i did not get your fraction clearly, can u write with mathematical symbols ?

OpenStudy (anonymous):

28/12x=802/12 divide by 28/12 you get x=802/12 times 12/28 x=802/28

OpenStudy (anonymous):

forget it thanks anyway i dont know how to use this site and i dont know how to get you to understand and i have no idea how to write them on here